a rectangle has sides of length 7 and 12, if the perimeter of a similar rectangle is 95. Find the length of its shorter side
Answer by edjones(8007)

2(7+12)=38 perimeter
95/38=5/2
7 * 5/2 = 35/2 = 17.5 length of the shorter side.
.
Ed